OpenStudy (anonymous):

A distance of one centimeter is the same as A~100 m B~ 10 m C~ .01 m D~1000 m

OpenStudy (danjs):

This is what you have to remember... to convert to the base Length unit m \[\frac{ 10^{-2} ~m}{ 1~cm }\]

OpenStudy (danjs):

= 1

OpenStudy (danjs):

A Centi-meter is 2 orders of magnitude smaller than a meter.. 10^-2 \[\frac{ 1~cm }{ 1 }*\frac{ 10^{-2}~m }{ 1~cm } = \frac{ 10^{-2} ~m }{ 1 }\]

OpenStudy (danjs):

If you have to convert units, first go to the base unit, then convert to the one you want, at least that is the way i do it (least to memorize)... good luck

OpenStudy (danjs):